# Causality Guided Representation Learning for Cross-Style Hate Speech Detection

### v1

@@ Abstract
The proliferation of online hate speech poses a significant threat to the harmony of the web. While explicit hate is easily recognized through overt slurs, implicit hate speech is often conveyed through sarcasm, irony, stereotypes, or coded language—making it harder to detect. Existing hate speech detection models, which predominantly rely on surface-level linguistic cues, fail to generalize effectively across diverse stylistic variations. Moreover, hate speech spread on different platforms often targets distinct groups and adopts unique styles, potentially inducing spurious correlations between them and labels, further challenging current detection approaches. Motivated by these observations, we hypothesize that the generation of hate speech can be modeled as a causal graph involving key factors: contextual environment, creator motivation, target, and style. Guided by this graph, we propose CADET, a causal representation learning framework that disentangles hate speech into interpretable latent factors and then controls confounders, thereby isolating genuine hate intent from superficial linguistic cues. Furthermore, CADET allows counterfactual reasoning by intervening on style within the latent space, naturally guiding the model to robustly identify hate speech in varying forms. CADET demonstrates superior performance in comprehensive experiments, highlighting the potential of causal priors in advancing generalizable hate speech detection.
Introduction
The proliferation of online hate speech poses a significant threat to the harmony of the web due to its complex and ever-evolving nature (Castaño-Pulgarín et al. 2021; Malmasi and Zampieri 2017; Mathew et al. 2019). While considerable progress has been made in identifying explicit hate speech characterized by overt hostility, slurs, and direct attacks, the detection of implicit expressions of hate conveyed through sarcasm, irony, stereotype, or coded language remains a critical unresolved challenge (Schmidt and Wiegand 2017).
@@ Introduction
Figure 1: Distribution of hate speech styles and target groups across various social media platforms.
@@
Consider the examples: (i) “F****** women asking for equal rights here. Why don’t you go whine and w**** around you floozy sl**.” and (ii)“I am so tired of doing everything on my own.. Need a woman so that chores can be taken care of…”. Both express misogynistic sentiments, but while the explicit version uses direct derogatory language that is easily flagged by conventional detection systems, the implicit version conveys the same prejudice through innuendo and rhetorical framing that is much harder to detect algorithmically. Further analysis of current hate speech detection systems reveals a persistent pattern: models trained predominantly on explicit hate speech data perform poorly when confronted with implicit expressions (Ocampo, Cabrio, and Villata 2023; Waseem et al. 2017). This limitation arises because these models often rely heavily on surface-level patterns such as specific slurs, aggressive syntax, or overtly hostile language that are typically absent in implicit hate speech, rather than the underlying hate intent that is shared across hate speech in various styles.
Moreover, hate speech disseminated on various platforms often targets distinct groups and exhibits unique stylistic variations, as demonstrated in Figure 1. This potentially induces spurious correlations between such factors and hate labels, which poses additional challenges for existing detection models. For instance, certain platforms might exhibit higher rates of explicit hate speech due to lax moderation, leading models to incorrectly associate platform-specific linguistic markers with hatefulness.
Motivated by these observations, we hypothesize that the generation of hate speech can be modeled as a causal graph involving key factors: contextual environment (e.g., platform), creator motivation (e.g., hate intent), target, and style, where (i) contextual environment is a confounder, which influence the the motivation, target, and style. (ii) Motivation, target, and style jointly shape the content of an online post. (iii) The hate label is determined by the creator motivation and contextual environment.
@@
Guided by this graph, we propose CADET, a causal representation learning framework for cross-style hate speech detection. CADET first projects online posts into a latent space and disentangles them into causality-aligned, interpretable factors, isolating genuine hate intent from superficial linguistic cues. Then it leverages a confounder mitigation module to control the confounding effect caused by the contextual environment. Furthermore, CADET allows counterfactual reasoning by intervening on style within the latent space, naturally guiding the model to robustly identify hate speech in varying forms.
@@
    We conduct extensive experiments on multiple real-world hate speech datasets, demonstrating that CADET significantly outperforms state-of-the-art baselines in cross-style generalization, providing robust, interpretable, and causally informed insights into hate speech characteristics.
Related Work
Hate Speech Detection
Early hate speech detection systems largely relied on lexicon-based filters, leveraging curated lists of slurs or offensive terms (Tanev 2024; Gitari et al. 2015). These methods delivered high precision for overt insults but performed poorly in recall, consistently missing hateful content conveyed through sarcasm, metaphor, or coded language. With the advent of labeled hate speech corpora, supervised machine learning approaches emerged, typically formulating detection as a binary text classification task. Early work employed logistic regression and SVMs on n-gram and TF–IDF features, sometimes augmented with user metadata (Talat and Hovy 2016; Davidson et al. 2017). While effective within-domain, such models often exploit dataset-specific artifacts, limiting generalization (Schmidt and Wiegand 2017; Fortuna and Nunes 2018). The transition to deep learning brought CNNs, LSTMs, and later transformer-based models such as BERT, which captured richer semantic context and established new benchmarks (Malik et al. 2024; Caselli et al. 2021; Mathew et al. 2021). Models like HateBERT (Caselli et al. 2021) retrained transformers on hate-centric corpora, and explanation-aware models like HateXplain (Mathew et al. 2021) further improved transparency. Despite progress, state-of-the-art models still struggle with implicit hate, such as dog-whistles or stereotypes, which lack explicit signals. Implicit hate evades both keyword filters and standard classifiers. Recent benchmarks explicitly highlight this gap: Latent Hatred (ElSherief et al. 2021) focuses on implicit hate, showing that standard models falter on such data. Other works examine coded hate, where benign-sounding terms (e.g., “skypes”) carry hidden meanings (Saleh, Alhothali, and Moria 2023), again demonstrating the need for contextual and connotative understanding. Datasets like HateCheck (Röttger et al. 2021) and DynaHate (Vidgen et al. 2020) stress-test models on adversarial and style-shifted examples, underscoring that evolving language and subtlety in hateful expression remain a moving target. These findings collectively show that a significant gap remains in handling implicit, subtle, or context-dependent hate. This motivates our work: CADET is designed to bridge this gap by focusing on the underlying intent, motivation, and context—beyond surface linguistic cues—enabling more robust and generalizable hate speech detection.
Domain Generalization in Hate Speech Detection
@@
Hate speech manifests across diverse online platforms and in varying styles, raising the challenge of domain generalization—whether models trained in one context can robustly transfer to others (Corazza et al. 2019; Phadke and Mitra 2020; Sheth et al. 2023). Studies consistently show that hate speech classifiers trained on one dataset or platform perform poorly when deployed elsewhere, due to reliance on domain-specific artifacts (Arango, Pérez, and Poblete 2019; Zhang, Chen, and Yang 2023). For instance, models fixating on particular slurs or minority identifiers may miss hate speech expressed differently on new platforms (Vidgen et al. 2019). To address this, research has explored pooling datasets from multiple domains (Arango, Pérez, and Poblete 2019), encouraging models to learn features less tied to any single source. Domain-invariant representation learning, especially via adversarial training, is widely used: a domain discriminator encourages the model to ignore platform-specific signals (Ganin et al. 2016). This has been adapted for hate speech with frameworks like MultiFOLD (Arango et al. 2024), which applies curriculum learning and adversarial losses to achieve cross-domain adaptation. Some works seek “universal” cues—such as sentiment or aggression—that might generalize, but such hand-engineered features risk missing critical nuance (Wiegand, Ruppenhofer, and Kleinbauer 2019; Sarwar and Murdock 2022). A particularly hard case is style shift: explicit versus implicit hate. Study demonstrates that transformer embeddings separate explicit hate well, but implicit hate clusters with benign content, creating a style gap (Ocampo, Cabrio, and Villata 2023). Their solution links implicit and explicit expressions targeting the same group, but full cross-style transfer remains elusive. Our work directly addresses these shortcomings. CADET explicitly disentangles style as a latent factor, leveraging counterfactual data augmentation: we generate and train on explicit/implicit variants of the same content, enforcing style-invariant predictions.
Casual Representation Learning for Hate Speech Detection
@@
Causal representation learning aims to discover latent variables that correspond to the true generative mechanisms behind data, rather than incidental correlations (Schölkopf et al. 2021; Ahuja et al. 2023; Brehmer et al. 2022). This is especially desirable in hate speech detection, where models should attend to harmful intent or prejudice rather than superficial cues tied to specific words or platforms (Sheth et al. 2024b). Recent work has applied causal reasoning to mitigate model bias: recent efforts attempt to construct a causal graph to identify how spurious correlations (e.g., stereotype-linked co-occurrences) drive false positives, using counterfactual data augmentation and multi-task learning to weaken such biases (Zhang, Chen, and Yang 2023). Other approaches generate synthetic examples by altering identity, style, or target in hateful posts, encouraging invariance to non-causal factors(Sen et al. 2022; Tiwari, Yuan, and Zhang 2022; Pamungkas and Patti 2019). A leading line of research seeks to disentangle causal factors in the representation space. The CATCH framework (Sheth et al. 2024b) models hate target as a confounding factor linked to platform distribution. By separating target-specific and invariant components, CATCH improves cross-platform transfer. However, most prior work only considers one or two latent factors, such as target or platform, which is not necessary for detecting deceptive implicit hatch speech. Our proposed CADET advances this direction by proposing a fine-grained causal model with four key latent factors: user motivation, target identity, hate style, and platform context. Our multi-part variational encoder disentangles these factors, with orthogonality and adversarial losses ensuring independence, bridging the gap between implicit and explicit hate speech detection.
@@
Figure 2: Causal graph of hate-speech generation. An unobserved contextual environment (U) shapes creator motivation (M), target (T), and style (S). M, T, and S jointly produce the post (X). The hate label (Y) is determined by M and U.
The Hypothesis: Hate Speech Generation
@@
To move beyond surface-level pattern recognition, it is essential to disentangle the causal properties that render content hateful from the spurious factors (especially stylistic variations) that determine how such hate is conveyed.
Motivated by this perspective, we introduce a causal graph that formalizes the generative process of hate speech as showcased in Figure 2. The graph captures the following components: (i) Creator motivation (M), denoting the latent drivers that compel individuals to produce hateful content, including intent, emotional state, perceived grievances, and anticipated impact. (ii) Target (T), representing the individual or group against whom the hateful expression is directed, typically identified by protected attributes such as race, religion, gender, or sexual orientation. (iii) Style (S), reflecting the expressive mechanisms that shape how hateful intent manifests in text, with particular emphasis on the distinction between explicit and implicit expression. (iv) Contextual environment (U), an unobserved confounder encompassing both platform-specific dynamics (e.g., moderation policies, user demographics, community norms) and broader sociopolitical conditions (e.g., current events, political climate, prevailing ideologies). (v) Post (X), the observable text emerging from the interaction of M, T, and S. (vi) Label (Y), the binary categorization of the content as hateful or non-hateful.
This causal graph highlights both the core challenges and actionable insights for cross-style hate speech detection. First, unobserved contextual confounders create spurious correlations between stylistic choices (S) and hate labels (Y). For example, platforms with strict moderation policies often pressure users to adopt more implicit forms of expression, thereby inducing misleading associations between platform-specific linguistic markers and hateful content. These same confounders also shape labeling practices, resulting in inconsistent standards across platforms and contexts. Mitigating such confounding effects is therefore essential for developing models that generalize reliably. Second, creator motivation (M) is the true causal determinant of whether content is hateful, whereas style (S) and target (T) only influence how hate is articulated. That is, the same hateful intent can manifest explicitly or implicitly, and can be directed toward different groups, without altering the underlying hatefulness. This perspective suggests that detection systems should prioritize capturing the invariant causal factor (M) over the variable stylistic or target-specific features. Finally, the graph motivates a principled path toward cross-style generalization through counterfactual reasoning. By intervening on the style (S) node while holding creator motivation (M) and target (T) fixed, we can construct counterfactual examples that express the same hateful intent in alternative stylistic forms. Training models to treat both original and counterfactual samples as equally hateful encourages representations that emphasize invariant causal signals rather than superficial stylistic cues.
Figure 3: Framework of the proposed method. CADET firstly decomposes the factor in the latent space guided by the causal graph. Then, it mitigates the confounding effect of the platform, where various constraints are applied to enable the meaning representation learning. various. Finally, latent counterfactual reasoning is employed to learn style-invariant features.
The Proposed CADET
@@
In this section, we formally define the problem (Section Problem Definition) and present CADET, a causality-guided representation learning framework for cross-style hate speech detection. CADET comprises three core components: (i) a causally-aligned disentanglement architecture, which separates generative factors in the latent space according to the causal graph (Section Causally-Aligned Disentanglement); (ii) a confounder mitigation mechanism, which reduces the impact of spurious correlations and alleviates confounding effects (Section Confounder Mitigation); and (iii) a counterfactual reasoning module, which promotes style-invariant representations by generating and leveraging counterfactual examples (Section Latent Counterfactual Reasoning).
Problem Definition
@@
$f^{*}=\arg\min_{f\in\mathcal{F}}\;\mathbb{E}_{(x,y)\sim\mathcal{D}_{s^{\prime}}}\big[\ell(f(x),y)\big],$ (1)
@@
Causally-Aligned Disentanglement
To enable the causality-guided representation disengagement, we first map online posts into a text space.
Text Encoding

### v2

@@ Abstract.
The proliferation of online hate speech poses a significant threat to the harmony of the web. While explicit hate is easily recognized through overt slurs, implicit hate speech is often conveyed through sarcasm, irony, stereotypes, or coded language—making it harder to detect. Existing hate speech detection models, which predominantly rely on surface-level linguistic cues, fail to generalize effectively across diverse stylistic variations. Moreover, hate speech spread on different platforms often targets distinct groups and adopts unique styles, potentially inducing spurious correlations between them and labels, further challenging current detection approaches. Motivated by these observations, we hypothesize that the generation of hate speech can be modeled as a causal graph involving key factors: contextual environment, creator motivation, target, and style. Guided by this graph, we propose CADET, a causal representation learning framework that disentangles hate speech into interpretable latent factors and then controls confounders, thereby isolating genuine hate intent from superficial linguistic cues. Furthermore, CADET allows counterfactual reasoning by intervening on style within the latent space, naturally guiding the model to robustly identify hate speech in varying forms. CADET demonstrates superior performance in comprehensive experiments, highlighting the potential of causal priors in advancing generalizable hate speech detection. [^1] Code: https://github.com/Shu-Wan/cadet
Hate Speech Detection; Causal Representation Learning; Domain Generalization; Counterfactual Reasoning
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792669isbn: 979-8-4007-2307-0/2026/04ccs: Computing methodologies Causal reasoning and diagnostics
@@
Figure 1. Distribution of hate speech styles and target groups across various social media platforms.
1. Introduction
The proliferation of online hate speech poses a significant threat to the harmony of the web due to its complex and ever-evolving nature (Castaño-Pulgarín et al., 2021; Malmasi and Zampieri, 2017; Mathew et al., 2019). While considerable progress has been made in identifying explicit hate speech characterized by overt hostility, slurs, and direct attacks, the detection of implicit expressions of hate conveyed through sarcasm, irony, stereotype, or coded language remains a critical unresolved challenge (Schmidt and Wiegand, 2017).
@@ 1. Introduction
Consider the examples: (i) “F****** women asking for equal rights here. Why don’t you go whine and w**** around you floozy sl**.” and (ii) “I am so tired of doing everything on my own.. Need a woman so that chores can be taken care of…”. Both express misogynistic sentiments, but while the explicit version uses direct derogatory language that is easily flagged by conventional detection systems, the implicit version conveys the same prejudice through innuendo and rhetorical framing that is much harder to detect algorithmically.
Further analysis of current hate speech detection systems reveals a persistent pattern: models trained predominantly on explicit hate speech data perform poorly when confronted with implicit expressions (Ocampo et al., 2023a; Waseem et al., 2017). This limitation arises because these models often rely heavily on surface-level patterns such as specific slurs, aggressive syntax, or overtly hostile language that are typically absent in implicit hate speech, rather than the underlying hate intent that is shared across hate speech in various styles. Moreover, hate speech disseminated on various platforms often targets distinct groups and exhibits unique stylistic variations, as shown in Figure 1. This can induce spurious correlations between these factors and hate labels, posing additional challenges for existing detection models.
@@
Motivated by these observations, we hypothesize that the generation of hate speech can be modeled as a causal graph involving key factors: contextual environment (e.g., platform), creator motivation (e.g., hate intent), target, and style, where (i) contextual environment is a confounder, which influences the motivation, target, and style. (ii) Motivation, target, and style jointly shape the content of an online post. (iii) The hate label is determined by the creator’s motivation and contextual environment. Guided by this graph, we propose CADET, a causal representation learning framework for cross-style hate speech detection. CADET first projects online posts into a latent space and disentangles them into causality-aligned interpretable factors, isolating genuine hate intent from superficial linguistic cues. Then it leverages a confounder mitigation module to control the confounding effect caused by the contextual environment. Furthermore, CADET allows counterfactual reasoning by intervening on style within the latent space, naturally guiding the model to robustly identify hate speech in varying forms.
@@
    We conduct extensive experiments on multiple real-world hate speech datasets, demonstrating that CADET significantly outperforms state-of-the-art baselines in cross-style hate speech detection, providing robust, interpretable, and causally informed insights into harmful online content moderation.
2. Related Work
2.1. Hate Speech Detection
Early detection systems rely on lexicon-based filters (Tanev, 2024; Gitari et al., 2015). While precise for overt insults, these methods suffer from low recall, failing to capture sarcasm, metaphor, or coded language. The advent of labeled corpora shifts the focus to supervised classification; early approaches utilize logistic regression and SVMs with n-gram or TF–IDF features (Talat and Hovy, 2016; Davidson et al., 2017). However, these models often overfit to dataset-specific artifacts, limiting cross-domain generalization (Schmidt and Wiegand, 2017; Fortuna and Nunes, 2018). The transition to deep learning–spanning CNNs, LSTMs, and transformers–enables the capture of richer semantic context (Malik et al., 2024; Caselli et al., 2021; Mathew et al., 2021). Domain-adaptive pre-training (e.g., HateBERT (Caselli et al., 2021)) and explainable architectures (e.g., HateXplain (Mathew et al., 2021)) further advance the state-of-the-art. Despite these gains, models remain brittle against implicit hate. Recent datasets on latent hatred (ElSherief et al., 2021), coded terms (Saleh et al., 2023), and adversarial shifts (Röttger et al., 2021; Vidgen et al., 2020) highlight significant gaps in handling subtle or context-dependent expression. CADET addresses this by prioritizing underlying intent and context over surface-level linguistics, enabling robust detection of implicit hate.
2.2. Domain Generalization in Hate Speech Detection
@@
Hate speech manifests across diverse online platforms and in varying styles, raising the challenge of domain generalization—whether models trained in one context can robustly transfer to others (Corazza et al., 2019; Phadke and Mitra, 2020; Sheth et al., 2023). Studies consistently show that hate speech classifiers trained on one dataset or platform perform poorly when deployed elsewhere, due to reliance on domain-specific artifacts (Arango et al., 2019; Zhang et al., 2023). For instance, models fixating on particular slurs or minority identifiers may miss hate speech expressed differently on new platforms (Vidgen et al., 2019). To address this, research has explored pooling datasets from multiple domains (Arango et al., 2019), encouraging models to learn features less tied to any single source. Domain-invariant representation learning, especially via adversarial training, is widely used: a domain discriminator encourages the model to ignore platform-specific signals (Ganin et al., 2016). This has been adapted for hate speech with frameworks like MultiFOLD (Arango et al., 2024), which applies curriculum learning and adversarial losses to achieve cross-domain adaptation. Some works seek “universal” cues–such as sentiment or aggression–that might generalize, but such hand-engineered features risk missing critical nuance (Wiegand et al., 2019; Sarwar and Murdock, 2022). A particularly hard case is style shift: explicit versus implicit hate. The study demonstrates that transformer embeddings separate explicit hate well while clustering implicit hate with benign content, creating a style gap (Ocampo et al., 2023a). Their solution links implicit and explicit expressions targeting the same group, but full cross-style transfer remains elusive. Our work disentangles style as a latent factor and leverages counterfactual reasoning to generate explicit/implicit variants of the same content, thereby enforcing style-invariant predictions.
2.3. Causal Representation Learning for Hate Speech Detection
@@
Causal representation learning aims to discover latent variables that correspond to the true generative mechanisms behind data (Schölkopf et al., 2021; Ahuja et al., 2023; Brehmer et al., 2022). This is especially desirable in hate speech detection, where models should attend to harmful intent or prejudice rather than superficial cues tied to specific words or platforms (Sheth et al., 2024b). Recent work has applied causal reasoning to mitigate model bias: recent efforts attempt to construct a causal graph to identify how spurious correlations (e.g., stereotype-linked co-occurrences) drive false positives, using counterfactual data augmentation and multi-task learning to weaken such biases (Zhang et al., 2023). Other approaches generate synthetic examples by altering identity, style, or target in hateful posts, encouraging invariance to non-causal factors(Sen et al., 2022; Tiwari et al., 2022; Pamungkas and Patti, 2019). A leading line of research seeks to disentangle causal factors in the representation space. The CATCH framework (Sheth et al., 2024b) models hate targets as confounding factors linked to platform distribution. By separating target-specific and invariant components, CATCH improves cross-platform transfer. However, most prior work considers only one or two latent factors, such as target or platform, which is insufficient for detecting deceptive implicit hatch speech. Our proposed CADET advances this direction by proposing a fine-grained causal model with four key latent factors: creator motivation, target, style, and context. CADET disentangles these factors through orthogonal and adversarial losses, ensuring independence and bridging the gap between implicit and explicit hate speech detection.
@@
Figure 2. Causal graph of hate-speech generation. Creator motivation (M), target (T), and style (S) are influenced by contextual environment (U). M, T, and S jointly determine the post (X). The hate label (Y) depends on genuine motivation M* and U. M* is independent of unobserved confounder U.
3. The Hypothesis: Hate Speech Generation
@@
To move beyond surface-level pattern recognition, we must separate the causal factors that render content hateful from the spurious factors that determine how such hate is conveyed.
To this end, we propose a causal graph that formalizes the generative process of hate speech as showcased in Figure 2. Post (X), representing the textual content produced by a creator, and the Label (Y), a binary indicator of whether the post is hateful. They are both accessible during training. The post X is generated by three latent factors: Creator motivation (M), capturing the writer’s intent for desired impact; Target (T), denoting the individual or group toward whom the post is directed; and Style (S), describing the linguistic expression of the content. These latent variables and the label are influenced by a shared, unobserved Contextual environment confounder (U). U represents platform-level factors such as moderation policies, demographics, as well as broader sociopolitical atmosphere. Differences in moderation strictness or audience tolerance can simultaneously shape a creator’s stylistic choices, target selection, and even how intent is shaped in practice, thereby inducing spurious correlations among M, T, S, and the label Y. To separate genuine intent from its context-dependent manifestation, we posit an additional latent variable M*, representing the creator’s true underlying motivation or intent, which is invariant to the contextual environment U and a direct cause of the label Y.
Unlike conventional causal formulations in domain generalization that assume domain shift affects either X or Y alone (Liu et al., 2021), our graph models domain shift as an unobserved confounder influencing both X and Y. This design reflects current online discourse, where content creators actively adapt their writing to the contextual environment U, often shifting from explicit to implicit or obfuscated expressions to evade moderation. As a result, classifiers trained predominantly on explicit-style corpora struggle to detect implicit hate in new domains. In our proposed graph, while the observed latents (M, T, S) are entangled with U, the true motivation M* is independent of the contextual environment. Hence, M* is a domain-invariant causal feature that is optimal and robust for out-of-distribution prediction (Wang and Veitch, 2022). Therefore, learning M* is the goal of CADET. Additionally, the graph also motivates a principled path toward cross-style generalization through counterfactual reasoning. By intervening on the style (S) node while holding creator motivation (M) and target (T) fixed, we can construct counterfactual examples that express the same hateful intent in alternative stylistic forms. Training models to treat both original and counterfactual samples as hateful encourages representations that emphasize invariant causal signals rather than superficial stylistic cues.
4. The Proposed CADET
@@
In this section, we formally define the problem (Section 4.1) and present CADET, a causality-guided representation learning framework for cross-style hate speech detection. CADET comprises three core components: (i) a causally-aligned disentanglement architecture, which separates generative factors in the latent space according to the causal graph (Section 4.2); (ii) a confounder mitigation mechanism, which reduces the impact of spurious correlations and alleviates confounding effects (Section 4.3); and (iii) a counterfactual reasoning module, which promotes style-invariant representations by generating and leveraging counterfactual examples (Section 4.4).
Figure 3. Framework of the proposed method. CADET firstly decomposes the factor in the latent space guided by the causal graph. Then, it mitigates the confounding effect of the platform, where various constraints are applied to enable the meaning representation learning. various. Finally, latent counterfactual reasoning is employed to learn style-invariant features.
4.1. Problem Definition
@@
(1) $f^{*}=\arg\min_{f\in\mathcal{F}}\;\mathbb{E}_{(x,y)\sim\mathcal{D}_{s^{\prime}}}\big[\ell(f(x),y)\big],$
@@
4.2. Causally-Aligned Disentanglement
To enable the causality-guided representation disengagement, we first map online posts into a text latent space.
4.2.1. Text Encoding